### v1

# FindingDory: A Benchmark to Evaluate Memory in Embodied Agents

Abstract
Large vision-language models have recently demonstrated impressive performance in planning and control tasks, driving interest in their application to real-world robotics. However, deploying these models for reasoning in embodied contexts is limited by their ability to incorporate long-term experience collected across multiple days and represented by vast collections of images. Current VLMs typically struggle to process more than a few hundred images concurrently, highlighting the need for more efficient mechanisms to handle long-term memory in embodied settings. To effectively evaluate these models for long-horizon control, a benchmark must specifically target scenarios where memory is crucial for success. Existing long-video QA benchmarks overlook embodied challenges like object manipulation and navigation, which demand low-level skills and fine-grained reasoning over past interactions. Moreover, effective memory integration in embodied agents involves both recalling relevant historical information and executing actions based on that information, making it essential to study these aspects together rather than in isolation. In this work, we introduce a new benchmark for long-range embodied tasks in the Habitat simulator. This benchmark evaluates memory-based capabilities across 60 tasks requiring sustained engagement and contextual awareness in an environment. The tasks can also be procedurally extended to longer and more challenging versions, enabling scalable evaluation of memory and reasoning. We also present baselines that integrate state-of-the-art VLMs with low level navigation policies, assessing their performance on these memory-intensive tasks and highlight areas for improvement. Our dataset and code will be made available here.
Figure 1: FindingDory: We introduce a benchmark designed to evaluate an agent’s memory capabilities by constructing scenarios that necessitate reasoning over previously collected experiences, including those involving various interactions. In these scenarios, an agent is provided with logs from prior interactions in indoor environments and tasked with completing navigation, pick, and place instructions. The figure illustrates a sample episode: first, a privileged agent, with full access to the environment, executes a set of randomly sampled navigation and manipulation tasks. Subsequently, the memory-evaluating agent is introduced into the same environment and tasked with completing instructions that require leveraging the logged experiences of the privileged agent to determine where to navigate or which object to retrieve.
**footnotetext: Equal Contribution
1 Introduction
Memory is a fundamental capability for intelligent agents that allows them to recall past experiences, adapt to dynamic environments, and make informed decisions over extended timescales. In humans and animals, it plays a crucial role in navigation, reasoning, and goal-directed behavior. As we strive to develop more capable embodied systems, equipping them with robust memory mechanisms is essential — particularly in settings where agents must process high-dimensional multimodal inputs to interact with their surroundings. The ability to store and retrieve relevant information is the key to unlocking sophisticated behaviors, from household assistance to autonomous exploration.
Recent advances in Vision-Language Models (VLMs) have significantly improved high-level reasoning and planning for embodied tasks. These models leverage large-scale multimodal training to exhibit strong scene understanding and action-guided perception. However, most existing VLM applications focus on short-term or static tasks, such as image captioning or Visual Question Answering (VQA), which do not require sustained memory usage. Extending these models to long-horizon embodied control introduces new challenges: unlike VQA, which often involves processing a single image or short video clip, embodied agents must integrate observations over time, recall past experiences, and act accordingly within their environment.
While recent efforts have sought to extend long-context understanding in VLMs—primarily through video-based QA or long-document comprehension [36, 44, 21]—these approaches fail to fully capture the complexities of long-term memory in decision-making. Many existing QA benchmarks either rely on multiple-choice formats, which enable guessing, or require extensive human annotation to curate question-answer pairs that require models to attend to long video sequences. In contrast, we propose embodied memory tasks—such as “navigating to a soft toy that you did not rearrange yesterday” or “find the second object picked up earlier”—that demand fine-grained recall and multi-hop reasoning over past observations. These tasks provide a more rigorous test of memory capabilities, requiring agents to track and respond to environmental changes rather than passively perceive static information.
In this paper, we introduce FindingDory, a benchmark designed to evaluate long-horizon memory in highly photorealistic simulated environments. Our benchmark consists of 60 diverse navigation tasks, which we categorize along various axes of memory requirements, that require both long-range temporal and spatial reasoning while discouraging heuristic shortcuts or random guessing. To ensure that performance is directly tied to memory utilization rather than simple perception, we curate scenarios where agents must recall past interactions to succeed. Additionally, our benchmark features dynamic environments - where agent interactions modify the scene - making it essential to reason over changing contexts. Furthermore, FindingDory is procedurally extensible, allowing tasks to scale in complexity as vision-language models improve, ensuring continued relevance as embodied AI progresses.
Unlike standard QA datasets that rely on subjective human annotations—such as manually selected keyframes or hand-crafted question-answer pairs—our benchmark leverages photorealistic simulation to enable automatic, objective evaluation of memory-based navigation. These tasks go beyond static recall since they require agents to perform complex spatio-temporal reasoning over past interactions. For instance, tasks such as “navigate to any object that you interacted with yesterday” demand not only accurate memory retrieval, but also strategic decision-making to identify and navigate to the nearest valid target from where the agent is currently positioned. In FindingDory, we introduce metrics that quantify task completion efficiency and not just successful recall. Our key contributions are:
  1. 1.
    A benchmark for evaluating long-horizon memory in embodied decision-making, featuring 60 diverse navigation tasks in highly realistic indoor environments.
  2. 2.
    A comprehensive evaluation of VLM based high-level policies combined with low-level navigation policies, analyzing their performance and limitations in memory-intensive tasks.
  3. 3.
    A systematic and extensible evaluation framework with metrics that enable future advancements in memory-efficient embodied agents.
2 Related Work
Table 1: Comparison of FindingDory with popular memory-focused embodied AI benchmarks.
Benchmark Photorealism Semantic Reasoning Task Specification Template Categories # Resulting Tasks Memory Length ($\pi^{*}$) Isolates Memory
MemoryMaze [27] One-hot $1$ 500-1000 steps
MemoryGym [28] One-hot $3$ 128-512 steps (Extensible)
MultiON [42] One-hot $1$ 500 steps
SIF [25] Language $3$ $240$ (val/test) ✗ (Uses semantic map)
OpenEQA (Active) [23] Language $1600$ 75 steps
GoatBench [16] Language + Image $1800-3600$ (val) 500 steps (Static Env)
Excalibur [52] Language $11$ 21 Requires exploration
Ours Language $11$ $\approx 6000$ (val) 500-3500 steps (Extensible)
In this section, we review related work on long-video question answering and embodied AI benchmarks that incorporate historical experience in their problem formulation. We defer the discussion of different approaches for augmenting policies with memory-like mechanisms to Section C.1.
2.1 Video QA benchmarks
Recent analyses of video QA benchmarks [24, 7] show that many tasks can be solved by attending to key frames, not full video sequences, suggesting that current benchmarks may favor keyframe selection over evaluating long-term memory. Long-context evaluations for foundation models also focus on “needle-in-a-haystack” tasks, evaluating retrieval of specific details rather than integrating information across many segments [50]. While recent video benchmarks feature more challenging questions [24, 19, 41, 11, 36, 35, 18, 30, 47, 51, 23], they adopt multiple-choice QA formats which are not suitable for evaluating fine-grained decision-making over long histories. They may also allow models to succeed through guessing and are not extensible since they require substantial human annotation effort.
2.2 Embodied AI benchmarks
Various benchmarks have been proposed to evaluate multi-modal language models in embodied settings where agents must leverage their experience in environments to improve future decision-making [42]. Many focus on embodied question answering (EQA), such as HM-EQA [31], which requires active exploration, and Open-EQA [23], which tests VLMs in a passive setting with pre-collected episode histories. Notably, Open-EQA found that a blind LLM baseline—one that ignores visual input—still performed well, suggesting that multiple-choice evaluation can introduce biases and reduce the need for genuine memory recall. Other benchmarks study the situated instruction following setting where agents must gather information to interpret ambiguous instructions [32, 20, 25]. However, many allow agents to sidestep memory challenges by precomputing relevant objects and locations based on recent history, effectively turning potentially memory-intensive tasks into structured instruction-following problems. Memory-Maze and Memory-Gym [27, 28] focus on partially observable decision-making with history-based state modeling, but their simplistic visuals and small state spaces limit their relevance to complex real-world memory settings.
In Table 1, we compare our proposed benchmark to the most related benchmarks within embodied AI and RL research with respect to the following desired properties:
  1. 1.
    Photorealism: Measures the visual complexity of environments, affecting the meaningful evaluation of foundation models trained on real-world data.
  2. 2.
    Scope of semantic reasoning needed over memory: Distinguishes between benchmarks that require agents to recall diverse semantic information versus a narrow set of attributes such as object locations.
  3. 3.
    Task Specification: Evaluates whether tasks are defined using flexible language-based descriptions, enabling richer memory challenges.
  4. 4.
    Memory horizon length: Adapting the definition from [26], this measures the temporal scope over which memory must be maintained for optimal task performance. Unlike static benchmarks requiring localized frame recall, dynamic environments demand reasoning over events distributed across extended observation sequences. This metric becomes difficult to quantify when tasks require substantial exploration phases.
  5. 5.
    Memory isolation: Whether benchmark performance isolates memory evaluation from confounding factors, particularly exploration requirements, which represent orthogonal challenges to memory recall and reasoning.
Our benchmark isolates memory evaluation by avoiding confounding related to exploration and focuses on careful and extensible task curation with automated metrics, requiring agents to recall and reason over long-horizon environmental observations.
3 FindingDory
In this section, we describe the design and construction of our embodied memory benchmark, FindingDory. We first outline the benchmark structure, then detail the constituent tasks and simulation setup.
3.1 Benchmark Overview
The goal of FindingDory is to evaluate an agent’s ability to retain and utilize past experiences to efficiently complete future tasks. Unlike standard navigation or perception-based evaluations, our benchmark explicitly tests memory by structuring tasks around past interactions rather than only immediate observations. To achieve this, we design a two-phase setup consisting of an experience collection phase, where scripted oracle agents autonomously gather experiences in the environment, followed by an interaction phase, where they are evaluated on their ability to recall and act upon that history. This decomposition allows for a controlled and standardized evaluation of memory, ensuring that the agent’s ability to remember past interactions is isolated from its ability to explore effectively.
The experience collection phase introduces dynamism into the environment by having the agent interact with and rearrange different objects. This ensures that the history contains meaningful state changes, allowing us to construct tasks that require spatio-temporal reasoning. Using oracle information from the simulator, the agent picks up objects from one location and places them onto designated receptacles (e.g., a table, chair, or bed). Episodes vary in length, spanning 400 to 3500 frames, with 2 to 11 object pick-and-place interactions (further details in Sec. B.2).
Once the experience collection phase ends, the interaction phase begins. Here, a task is sampled from a predefined set of templated instructions, requiring the agent to recall and act upon the state of the environment from the previous phase. The agent is provided with access to its full history, including recorded images, pose, and past actions, and is tested on its ability to integrate long-term memory into decision-making.
Due to the photorealistic visuals and object diversity of the Habitat simulator, FindingDory provides a practical and transferable setting for evaluating large-scale vision-language models. Since many VLMs are pretrained on real-world images, they can readily interpret frames from this environment—enabling realistic memory evaluations in simulation, where long-horizon, multimodal interaction data can be collected at scale. This makes FindingDory especially suitable for studying how current foundation models integrate memory into decision-making.
Table 2: Task Categories with Example Prompts and their Challenges. See Appendix F for detailed examples.
{tblr}
width = colspec = X[0.04,l,m]X[0.9,l,m]X[3.0,l,m]X[2.1,l,m], hline1,2,13, row1 = font=, bg=lightgray!30, roweven = bg=lightgray!10, cell12,3,4 = font=, c, column1 = colsep=6pt, column2 = colsep=1pt, column3 = colsep=1pt, column4 = colsep=1pt, & Task Category Example Instructions Memory Requirement
[…]
Interaction Navigate to any object that you did not interact with yesterday. Differentiate between interacted & uninteracted objects.
Conditional Interaction Navigate to a {receptacle_category} you picked an object from. Object categories, object-receptacle relationships.
Object Attribute Navigate back to a {target_color} colored object that you interacted with yesterday. Object attributes other than category.
Spatial Relationship Navigate to the object which you interacted with which is the farthest from your current location. Spatial relationships to objects.
Room Visitation Navigate to a room that you did not visit yesterday. Room boundaries and past interactions.
[…]
Time-Based Navigate to the object that you interacted with at {HH:MM} yesterday. Time of interaction & final position of objects.
Duration Tracking Navigate to the object which took the longest time to rearrange. Duration of interactions.
[…]
Ordered Revisitation Revisit all the objects you interacted with yesterday in specific order. Similar to unordered revisitation but with strict sequencing.
3.2 Tasks
Our benchmark is designed to evaluate a broad spectrum of memory challenges faced by embodied agents, moving beyond narrow tasks that could be solved using simple heuristics or non-generalizable strategies. We define several templated tasks that evaluate an agent’s ability to integrate and recall different forms of memory across spatial, temporal, and semantic dimensions. Some tasks require recalling specific objects or receptacles, tracking when (e.g., temporal ordering) and how (e.g., interaction type: picked vs. placed). Others demand awareness of room-level context (e.g., rooms visited, farthest location) or semantic attributes (e.g., shape, color, material, functionality). We also introduce time-usage memory tasks, such as identifying the room that took the longest to explore or the object with the longest rearrangement duration. Task instructions are generated by randomly sampling scene-specific entities, and the full list of 60 templates is shown in Table LABEL:tab:all_tasks.
A smaller subset of attribute-based evaluations is also included to assess whether agents can flexibly reason over visual and semantic memory. These tasks involve prompts that refer to objects indirectly via properties observed during prior interaction (e.g., “go to the striped object you interacted with previously”). This setup helps differentiate methods capable of recalling fine-grained visual details from those relying solely on static representations, such as semantic maps [8] or textual summaries [13], which often struggle to flexibly reconstruct such information.
Single vs. Multiple Goals. The majority of tasks in FindingDory involve single-goal instructions, where the agent must reason over its experience to identify a target and navigate to it. A smaller subset contains multi-goal tasks that require reaching multiple destinations, sometimes in a specific order. Single-goal tasks can be equally challenging, often requiring agents to recall what was not interacted with or to trace specific object movements through rearrangement states—demanding more than simple recall from static frames.
Extensibility through Procedural Generation. Our task suite is procedurally extensible, allowing difficulty to scale with changes in the experience collection phase. Increasing the number of pick-and-place interactions, for instance, expands the temporal reasoning space and creates more challenging memory dependencies.
Task Categorization. We group tasks into interpretable categories to analyze performance trends. Table 2 presents these categories along with representative instructions and the memory challenges they embody. To our knowledge, FindingDory offers the most visually complex and extensible suite of embodied memory tasks to date, covering a wide range of interaction-driven and long-horizon reasoning scenarios.
3.3 Implementation details
Simulation Setup and Dataset Composition. We build FindingDory using the Habitat simulator [33, 37] and scenes from the Habitat Synthetic Scenes Dataset (HSSD)[15]. Our benchmark uses 107 scenes from the training split and 30 from the validation split, resulting in 1,478 training and 100 validation episodes. We populate these scenes with the same object sets used in OVMM[48], combining assets from AI2Thor [17], Amazon-Berkeley Objects [9], Google Scanned Objects [10], and HSSD [15].
Across these episodes, we include 839/247 object instances spanning 84/72 distinct categories in the train/val splits, respectively. Objects are rearranged during the experience collection phase to induce temporally grounded memory challenges. Additionally, agents interact with 17/16 distinct receptacle categories in the train/val splits (e.g., tables, chairs, beds). To support attribute-based tasks (see Table 2), we generate semantic descriptions of object instances using GPT-4o. We prompt the model with multi-view images of each object to extract attributes such as shape, color, material, and functionality. These descriptions are used to populate the object attribute task templates (5 total), which are limited to the validation set and undergo manual quality review.
In total, our benchmark includes 82,174 training and 5,876 validation task instances. Details on the procedural generation pipeline for episodes and instructions are provided in Appendix B.
Agent Setup. We adopt the Hello Robot Stretch embodiment [14] in our simulation. The agent is equipped with RGB and depth sensors (resolution $640\times 480$) and a GPS+Compass sensor. The agent can execute one of four discrete navigation actions: MOVE_FORWARD ($0.25$m), TURN_LEFT / TURN_RIGHT (by $10^{\circ}$) and a STOP action. During the experience collection phase, the agent can also manipulate its arm joints (extension, lift, yaw, pitch, roll) and activate a manipulation_mode for pick-and-place actions, which rotates the base ($90^{\circ}$) and directs the head camera toward the gripper.
PDDL-based Task Verification. Since we operate with a large and diverse set of instruction templates (see Table 2), we leverage the PDDL [2] specification system as instantiated in [38]. For each instruction template, we create a template goal condition that uses various predicate expressions such as is_robot_at_object(x, "OBJECT" or is_robot_at_receptacle(x, "RECEPTACLE" where x is some placeholder variable. These placeholder variables are sampled based on various constraints such as target_color or interaction_order that are defined through the template. Together, the placeholder constraints and goal conditions enable us to create scalable, automated verification routines for each task in FindingDory.
4 Experiments
We evaluate several baseline approaches on FindingDory and assess their ability to complete memory-intensive tasks using a combination of vision-language reasoning and navigation. Below, we describe the baseline architecture, implementation variants, and evaluation metrics.
Figure 2: FindingDory Task Performance: Proprietary VLMs such as Gemini-2.0-flash and GPT-4o fail to perform appreciably across most FindingDory task categories. Agents struggle significantly on multi-goal tasks where multiple subgoals are to be achieved sequentially based on the interaction videos collected during Phase 1 of the task (sometimes in a specific order). The Qwen-SFT baseline (see Sec. D.5) which is trained to predict ground-truth frame indices performs best but still reaches only $\approx 50\%$ on the high-level tasks. This highlights a considerable gap in the spatio-temporal reasoning capabilities of frontier models.
4.1 Evaluated Approaches
To establish strong baselines, we adopt a hierarchical architecture (Fig. 11) comprising two modules:
(1) a high-level goal selection module, which processes the interaction history to select goal frames, and
(2) a low-level navigation policy, which executes actions to reach the selected frame.
This architecture enables reasoning over long-horizon multimodal experience while delegating fine-grained control to a learned policy. Inspired by prior work [45], we avoid directly predicting actions using VLMs due to their poor performance in continuous control, and instead restrict them to high-level goal prediction.
4.1.1 High-Level Goal Selection
Video LM Agent. This agent receives the full interaction video along with a task instruction. Each frame in the video is annotated with its index and the time of day at which it was captured. The agent is tasked with predicting the index of a frame from the video that corresponds to a viable goal state. Specifically, the predicted frame should represent a viewpoint from which, if the agent were to navigate back to the same camera position, it could successfully complete the task. Although the model outputs a single frame index, selecting the correct one often requires reasoning over multiple temporally scattered observations, including frames where the agent may not be actively interacting. Since different VLMs are capable of handling different numbers of frames in context, we subsample the frames passed to each model based on what empirically performs best. We evaluate both proprietary (Gemini-2.0-Flash [39], GPT-4o) and open-source (Qwen2.5-VL [5], Gemma-3 [40]) vision-language models capable of processing long video contexts. See Secs. D.1, D.2 and D.3 for more details.
Textual Memory Agent. Similar to [25, 4], we also consider an explicit, memory-building baseline which breaks down the video frames into chunks and uses a VLM to generate textual summaries for each video chunk. After “captioning” the entire video sequence, the VLM selects frame indices by reasoning over the generated textual descriptions. See Sec. D.4 for more details.
Supervised Fine-Tuning (SFT). We additionally explore a supervised fine-tuning approach where the high-level VLM is trained to predict goal frame indices, given the task prompt and interaction history. Since multiple frames may validly represent the goal (e.g., several consecutive frames showing an object being placed), we supervise the model on the full list of acceptable target frames. At test time, a single frame is sampled from the list of predictions and is considered correct if it matches any of these valid frames. This setup enables the model to learn task-specific cues from examples and serves as a stronger baseline than zero-shot prompting, particularly for tasks requiring subtle temporal grounding. We provide additional details in Sec. D.5.
4.1.2 Low-Level Navigation
To execute the high-level plan, we use a low-level policy that takes the selected goal frame and outputs discrete actions: MOVE-FORWARD, TURN-RIGHT, TURN-LEFT, and STOP. Our primary policy is the LSTM-based controller from OVRLv2 [46], trained to reach image goals from egocentric RGB-D observations (see Appendix E.1).
We also evaluate a mapping-based policy that deterministically navigates to the location corresponding to the goal frame using a global map, without relying on learned visual features (Appendix E.2).
Solvability. Since the action space of our high-level baseline is constrained to frames observed during the experience collection phase, there are tasks for which this is insufficient. For example, in the task “Navigate to a room that you did not visit yesterday,” all available waypoints correspond to previously visited locations, making it impossible to satisfy the instruction regardless of the selected frame. To establish an upper bound on task difficulty, we quantify the percentage of episodes that could theoretically be solved through optimal goal selection from interaction history, assuming perfect navigation (teleportation) to the selected waypoint. This upper bound is the dashed line (Oracle) in Fig. 2.
(a) Relaxed metrics for Objects.
(b) Relaxed metrics for Receptacles.
(c) HL-SR vs HL-SPL Comparison.
(d) Hierarchical policy results.
Figure 3: (a, b) Relaxed metrics for objects and receptacles tasks. We find that the models significantly improve in relaxed metrics for tasks involving receptacles while tasks targeting smaller objects do not exhibit similar improvements. (c) Success Rate vs SPL performance On tasks in which the VLM has the option to choose between multiple correct frames, we find that all models choose suboptimal frames sometimes resulting in the reduced HL-SPL metric. (d) Hierarchical policy results on FindingDory When Qwen is combined with an ImageNav or Mapping agent, the overall SR and SPL reduces significantly (seen in the LL-SR/LL-SPL metrics).
4.2 Evaluation Metrics
We benchmark agent performance on the FindingDory tasks using Low Level Success Rate (LL-SR), assessing whether the agent finds the right target entities, and Low Level Success-weighted-by-Path-Length (LL-SPL), evaluating if the agent selects and navigates to the optimal entity using the shortest possible path. Since we use a hierarchical baseline, we additionally introduce a High-Level Policy Success Rate (HL-SR), and High-Level Policy SPL (HL-SPL) which measures the accuracy and efficiency of the high-level policy in selecting the correct and closest goal frames for navigation. To further analyze failure modes, we also report two relaxed variants: Distance-to-Goal-Only SR (DTG-SR), which relaxes the requirement that the target object must be visible in the selected frame, and Semantic Coverage SR (SC-SR), which ignores spatial proximity. Additional metrics, success criteria, and threshold specifics are provided in Sec. B.3.
5 Results
In this section, we analyze the performance of the outlined baselines on FindingDory, presenting results grouped according to the categories described in Section 3.2.
State-of-the-Art VLMs struggle on the FindingDory benchmark. From  Fig. 2, we observe consistently low success rates on all task subsets across baselines, with GPT-4o achieving the highest performance at 27.3%, Gemini-2.0-Flash reaching 25.7%, while Qwen2.5-VL and Gemma-3 score in the 13–16% range. The text-based agent performs the worst overall, underscoring that textual memory representations alone are insufficient for disambiguating visual goals or reasoning over complex visual histories. Notably, the supervised fine-tuned (SFT) variant outperforms all frozen VLMs, achieving an average improvement of 25% across tasks, with considerable improvement on temporal and multi-goal tasks.
In single-goal tasks, all methods struggle particularly with the spatial relationship subset, with even GPT-4o achieving only 14.5% success. These tasks require reasoning about relative distances between entities and the agent—an ongoing challenge noted in prior work [47]. In contrast, room identification tasks yield the highest performance, likely due to the small number of distinct rooms and the lenient goal condition of reaching any location within the correct room. Performance on conditional interaction and object attribute tasks is comparable, suggesting that current VLMs handle both category-based queries (e.g., “apple”) and attribute-based descriptions (e.g., “red food”) with similar effectiveness.
Temporal tasks pose additional challenges to the existing VLMs, with models consistently failing to identify the correct order of interactions—requiring multi-hop reasoning. Duration tracking is similarly difficult as it requires inferring elapsed duration based on specific points in the video. In contrast, for the time-based tasks we observe that baselines perform better as these tasks only require single frame retrieval without additional reasoning.
(a) Comparison of different video lengths
(b) Reasons for ImageNav Policy Errors
Figure 4: (a) Model performance at different number of subsampled video frames. (b) Reasons for ImageNav Policy Errors.
Poor performance on multi-goal tasks. From Fig. 2, we observe that performance of each VLM baseline that we consider is close to zero on most multi-goal tasks. This suggests that VLMs find it hard to track and remember multiple object-receptacle interactions. For the multi-goal evaluations, we use a single-shot prompt and require the high level VLM agent to predict all target frame indices in one single response - primarily to avoid multiple inferences with long sequence lengths due to large number of frames. Qualitatively, we observe that the VLM generally fails to even predict the correct number of targets required to solve the multi-goal task. Our trained Qwen model performs slightly better, achieving close to 20% accuracy, but still falls significantly short of the achievable result of 99%.
VLMs can identify target receptacles but not precisely locate them spatially. We conduct an experiment to analyze the performance differences of VLMs between tasks involving objects versus receptacle entities. We extract a subset of FindingDory tasks, categorizing them based on whether the valid entities are objects or receptacles. From  Fig. 3(b), we observe that the relaxed Semantic Coverage-SR (SC-SR) metrics (see  Section B.3 for details) significantly increases for tasks involving receptacle entities. The marked increase in high-level success rates with relaxed distance threshold (DTG-SR) suggests that VLMs are generally able to identify frames from which large entities (receptacles) can be viewed in the image but cannot accurately identify specific frames which are spatially closest to the target entity. In contrast, tasks targeting smaller objects do not exhibit the same improvement (see  Fig. 3(a)), as fewer number of frames clearly show these small objects from farther distances. This highlights a significant limitation of current VLMs in constructing effective spatial representations to accurately localize target entities from the provided observation stream. We explicitly prompt high-level VLM agents to identify frames spatially closest to target entities, essential for task success (see Figs. 6 and 8 for example prompts).
VLMs struggle with identifying the closest entity that solves the task. In Fig. 3(c) we report the HL-SPL metric (Section 4.2) which captures the efficiency with which the high level VLM-based modules select goal frames. During the experience collection phase, the agent can interact with multiple object instances of the same category due to which a task such as “Navigate to a squeezer you interacted with” could have multiple valid entities that the agent can navigate to. We observe that for such tasks there is a wide gap between HL-SR and HL-SPL (upto $50\%$ in case of interaction-based tasks) for all model variants. This suggests that although the VLMs can correctly recognize a valid entity to solve the task, they still fail to perform a fine-grained spatial analysis of the multiple valid entities being rearranged dynamically in the video. Thus, they cannot spatially locate the valid entity nearest to its current position.
Policy decomposition causes performance degradation. To benchmark the hierarchical agent proposed in  Sec. 4.1, we experiment using Qwen2.5-VL-7B as the high-level VLM module and image-goal navigation policy and mapping policy as the low-level agents. The low-level policy activates only when the high-level VLM correctly selects the subgoal. As depicted in  Fig. 3(d), the Low-Level-SR (LL-SR) metric significantly drops across all FindingDory tasks for the image-goal navigation policy. We attribute this decline to a distribution shift in the visual goals provided to the navigation policy. During training, the policy encounters randomly sampled start and end goal poses, differing substantially from those selected by VLMs during evaluation. Successful VLM-selected frames typically correspond closely to pick place interaction routines, resulting in image goals predominantly containing background walls and partial views of receptacle surfaces. Consequently, this limited visual context hinders effective navigation. In comparison, the mapping based navigation policy performs better, having only a 25% relative drop in SR. The SPL of both the methods remain in the single digits. This emphasizes the importance of the FindingDory benchmark in revealing these gaps which motivates the development of more reliable memory and navigation strategies
Frame Subsampling. We analyze how the number of input frames affects high-level performance by evaluating agents on subsampled interaction videos of varying lengths (Fig. 4(a)). This helps assess whether models benefit from longer histories or simply overfit to sparse visual cues. Interestingly, we observe that frozen VLMs—despite being capable of accepting long contexts—do not improve with more frames. In fact, their performance often degrades at higher frame counts (minimal subsampling), suggesting that they struggle to extract relevant information from densely packed, unfiltered input. This highlights a key limitation: naïvely increasing context length does not help unless the model can effectively attend to localized events.
In contrast, the fine-tuned model (Qwen SFT) demonstrates clear gains when trained with longer videos, leveraging additional context to improve reasoning. This indicates that with appropriate supervision, models can move beyond shallow matching and utilize richer temporal signals from the interaction history.
Sanity checks. We investigate how a VLM would perform at short-horizon visual comprehension on images from the same episodes used for long-horizon evaluation in our benchmark. This tests if poor performance of the evaluated models on our long-horizon reasoning tasks is due to a domain or sim2real gap. We assess this by sampling frames around object interactions, asking models to list out detected objects, and check accuracy compared to the ground truth category via string matching – a conservative metric that misses correct predictions using synonyms or alternative phrasings. This still results in image sequences of length 50-150, but Gemini achieves an accuracy of 67% which is significantly higher than it’s average high level success rate across tasks in Fig. 2. This shows that despite clear, close-up views, models struggle to reason over them when temporal sequence comprehension is required.
Low level navigation policy failures. From  Fig. 4(b), we observe that the low-level image-goal navigation policy can fail due to invoking the termination (stop) action incorrectly or a navigation routine timeout (max budget of $1000$ steps including the data collection phase). To further understand this behavior, we analyze if the navigation policy actually looks at the target entity when the stop action is invoked. Interestingly, we find that in several instances the target entity is visible from a slightly larger distance threshold (2.0m) from the chosen task success threshold (1.0 m, in case of receptacles). This often happens when VLMs select goal frames in which the target entity is visible in the 2D image but not necessarily closest in 3D space, and thus the goal frame coordinates are slightly further away than the threshold distance.
6 Conclusion
We introduced FindingDory, a benchmark for evaluating long-horizon memory in embodied agents within a highly photorealistic indoor simulation. Our findings reveal limitations in VLM context scaling, particularly in handling long observation sequences required for memory-intensive tasks. FindingDory provides informative metrics that disentangle different aspects of memory reasoning, allows for procedural task generation of increasing difficulty, and fine-grained, efficiency-focused failure analysis across high-level semantic retrieval and low-level spatial reasoning. Its extensibility allows for longer temporal dependencies, supporting research on context scaling and memory efficiency in multimodal models. In our results, we observe that the supervised fine-tuning baseline performance improves consistently when trained with longer videos which suggests future work should explore finetuning VLMs with effective memory compression techniques for enhanced spatio-temporal reasoning.
Our experiments also highlight challenges with hierarchical policies that integrate long-context VLMs with low-level control. While high-level policies specify abstract goals—such as referencing a past frame in the interaction history—low-level navigation policies often struggle to reach the intended object precisely, exposing a gap between goal specification and reachability. Addressing this limitation requires better integration between high-level reasoning and low-level execution, either by improving hierarchical policies to leverage VLM guidance more effectively or by training end-to-end models that jointly learn memory retrieval and control.
Beyond evaluation, FindingDory can serve as a training platform for memory-dependent tasks, enabling long-context-aware models and enabling dataset extraction for supervised learning tasks such as VQA over long videos. By bridging the gap between VLM advancements and real-world memory demands, FindingDory facilitates systematic study of memory-augmented policy learning and long-term reasoning, driving progress in multimodal embodied AI.
Supplementary Material
Appendix A Limitations
While FindingDory provides a rigorous evaluation framework for long-horizon memory in embodied AI, it has certain limitations that highlight areas for future improvement.
First, our most competitive baseline, which integrates a VLM with a navigation policy, is constrained by its low-level controller. Current memory-guided navigation policies struggle to take the shortest paths in large environments after reasoning over long interaction histories. As a result, models perform poorly on spatial awareness metrics, making it difficult to assess their true memory capabilities beyond basic path efficiency. Future work should explore improved memory-driven navigation policies to better leverage long-term spatial reasoning.
Second, in the experience collection phase, our scripted oracle policy for object interactions (Magic Grasp) executes pick-and-place actions in an unnatural way—objects are abruptly transferred from the receptacle surface to the agent’s gripper without smooth motion. This may confuse VLMs, as the sudden change is visually subtle and may be overlooked if the model does not attend to the specific frames where it occurs. Developing more naturalistic interaction models could improve the realism and interpretability of memory traces.
Third, while we introduce distractor objects to discourage guessing, we do not explicitly control how prominently these objects appear in the experience collection phase. Currently, objects the agent interacts with are often viewed more frequently and closely compared to distractors, which tend to be featured less prominently in the agent’s field of view. Although some tasks mitigate this by explicitly referencing non-interacted objects, future iterations of the benchmark may need stronger distractions by ensuring more balanced exposure between distractors and interacted objects.
Finally, while the experience collection phase involves both navigation and manipulation, the current interaction phase tasks are all navigation-based, as memory reasoning is primarily required for locating objects rather than executing pick-and-place actions, which are Markovian once the goal is known. However, our benchmark is easily extensible to include manipulation-based memory tasks, which would further enrich the evaluation of long-term reasoning in embodied AI.
Despite these limitations, FindingDory establishes a strong foundation for studying long-horizon memory, offering extensible challenges that will evolve alongside advances in vision-language models and embodied learning.
Appendix B Benchmark Details
B.1 FindingDory Episode Creation
In this section, we give details about our episode creation pipeline that allows us to procedurally generate data for our benchmark.
Object-Receptacle Pairing. The first step in the episode creation involves generating valid spawning positions for various object-receptacle pairs by sampling entities from the respective split in consideration (train/val). The objective is to identify candidate objects that can be picked up from a start receptacle and placed stably on a goal receptacle. We use the pipeline proposed in [48] to procedurally generate various object placements on receptacle surfaces by running physics checks for stable placements. To make the generation procedure easier, we assume that each object is picked and placed only once during an episode. Additionally, we constrain the start receptacle category to be distinct from the goal receptacle category. We also ensure that for each candidate object category, a corresponding non-interacted object of the same category is spawned to serve as distractor entities during the interaction phase.  Figures 5(b) and 5(c) provide a distribution of distractor objects and receptacles in the val episodes. The number of sampled object-receptacle pairings is equal to the number of rearrangements in the particular episode and can be configured through an argument. In essence, this controls the distribution of video frame lengths in the FindingDory tasks(see  Figure 5(a)).
Object Placement Sequences. Once we have identified the set of candidate objects and the corresponding goal receptacles, we require an oracle policy through which we can “naturally” drop objects on the receptacle surfaces. In practice, one can achieve this by training a “place” skill through end-to-end reinforcement learning or use a heuristic agent that can place objects on receptacles but both frameworks do not guarantee perfect $100\%$ success which is crucial for solvability in FindingDory tasks. To address this, we construct an offline pipeline that produces the set of low-level “oracle” joint actions that lead to stably placing the object on the goal receptacle. We spawn the agent (with the object snapped in its end-effector) at the associated viewpoints of goal receptacles on the navigation mesh. From each viewpoint, we try to execute a heuristic, state-machine based placing policy which performs a set of predefined joint effector actions after identifying a “clutter-free” region in the receptacle pointcloud (using an onboard depth sensor). The output of this routine produces a set of low level joint actions and the corresponding receptacle viewpoint for each goal receptacle that was identified in the previous object-receptacle pairing stage. We reject episodes where the heuristic-based place policy cannot stably place the object from any of the associated goal receptacle viewpoints.
Task Creation. Once we have the final filtered set of episodes, we populate task instructions for each episode using yaml-based instruction templates as used in [38]. We employ the instruction templates to procedurally populate various placeholders such as object names, attributes and appearance orders that create specific task instructions grounded in the entities (objects and receptacles) present in the episode in an automated manner. We present the diverse instructions templates we use for each task category in Table LABEL:tab:all_tasks. In total, we create $82174/5876$ unique task instructions within the train/val split respectively.
(a) Distributions of frame counts and interaction counts across episodes.
(b) Number of objects per episode.
(c) Number of receptacles per episode.
Figure 5: Summary statistics of episode data.
B.2 Experience Collection Phase
We now outline the details of the experience collection phase grounded in the episodes created by the pipeline outlined in the previous section. The experience collection phase is designed to collect a “clean” and “natural” video sequence of an agent interacting in an home environment while interacting with various entities. We focus on object rearrangements as the core interaction routine followed by the embodied agent. We construct an OracleAgent agent that uses privileged information from the simulator instance to efficiently generate such interaction routines. We now detail the OracleAgent policy to generate such video sequences for a particular episode.
Nav-to-Pick. The agent begins by being spawned at a random location in the scene associated with a particular episode. We then query a ShortestPath function that leverages privileged simulator information to navigate to the object that has to be picked up. The OracleAgent follows the shortest path to the object and collects images as it navigates in the environment.
Pick. Once the OracleAgent reaches the object location, it executes a hardcoded pick policy sequence which results in the object being magically grasped by the OracleAgent. The pick policy uses the onboard depth sensor to orient itself with the object instance placed on the receptacle infront of the agent. Once the object is in clear view (using the ground truth instance mask from the simulator), the and effector extends outwards to perform a magic_grasp action. In cases when the object is not visible, we re-query the ShortestPath function to navigate to an alternate viewpoint associated with the receptacle that would lead to the object being clearly visible.
Nav-to-Place. After grasping the object, the OracleAgent navigates to the corresponding goal receptacle that was sampled during the episode creation phase. Specifically, the agent navigates to the specific viewpoint associated with the goal receptacle from which the pre-recorded placement action sequence is to be executed.
Place. Once the agent reaches the specific viewpoint, we re-play the end-effector action sequence that was found to lead to a stable object placement during episode creation. This marks the end of a single rearrangement interaction routine and the agent transitions to rearrange the next candidate object in the episode until no more objects are left.
Final Navigation. At the end of all interaction routines within an episode, the agent navigates to a predefined location to ensure that it does not remain in close proximity to the final interacted object or the final goal_receptacle. This ensures that tasks are not solvable by selecting only the last few frames in the video sequence. We ensure that the agent navigates to a distance of $3$ m away from the final interaction location.
The above OracleAgent routine enables us to procedurally create videos of the agent rearranging objects in the environment. Additionally, as the agent navigates the environment between the pick and place substeps, it collects additional frames detailing the scene environment which can be useful (or serve as distractor frames) when attempting to solve FindingDory tasks. We present an example experience collection trajectory and associated “magic” pick-place sequence in Figures 12 and 13.
B.3 Evaluations
In this section, we define the primary evaluation metrics employed to benchmark the performance of various baselines on the FindingDory tasks. Following [6], we focus on metrics that systematically quantify both the success and efficiency of the agent in solving specified instructions. Success focuses on accurately selecting frames from the interaction video that lead to task completion, while efficiency evaluates if the agent selects and navigates to the optimal frame that solves the task using the fewest low-level control actions. Notably, optimizing efficiency poses a significant challenge for current methods [29, 47].
High-Level Policy Success Rate (HL-SR). This metric measures the percentage of episodes in which the high-level agent correctly predicts frame indices required to solve the specified task. We define an episode as successful if the frame(s) predicted by the VLM satisfy the following criteria:
  • Distance-to-Goal: The agent is within a specified distance of the target entity upon reaching the predicted frame. We use thresholds of 2.0 m for objects and 0.1 m for receptacles. The larger threshold for objects accounts for the possibility of objects being positioned on elongated receptacles (e.g., couches), potentially affecting reachability.
  • Angle-to-Goal: The agent is oriented within a specified angular distance toward the target entity’s center. We set angle thresholds of $45^{\circ}$ for objects and $90^{\circ}$ for receptacles due to their typically larger dimensions.
  • Semantic Coverage: The semantic mask of the target entity covers at least $0.1\%$ of the pixels in the frame.
  • Room Region Check: For room visitation tasks, we verify if the agent is located within the designated target region.
These thresholds and criteria were qualitatively validated through manual inspection to ensure their appropriateness; During evaluation, these criteria are verified by teleporting the robot to the pose associated with the frame(s) selected by the high-level agent.
Low-Level Policy Success Rate (LL-SR). This metric measures the percentage of episodes where the low-level navigation agent successfully navigates to the target entity. The low-level agent is activated only if the high-level agent correctly predicts the frame indices, meaning LL-SR is evaluated conditionally upon HL-SR being true for each episode. The success criteria match those outlined for HL-SR.
Success-weighted-by-Path-Length. This metric evaluates task completion efficiency, computed per task episode as follows [6]:
$\text{SPL}_{i}=S_{i}\cdot\frac{l_{i}}{\max(p_{i},l_{i})}$
where $S_{i}$ denotes HL-SR/LL-SR for computing the HL-SPL/LL-SPL respectively, $l_{i}$ is the length of the shortest possible path to the closest successful goal frame (or shortest path to revisit all subgoals), and $p_{i}$ is the actual path length traveled by the agent.
Additionally, we introduce two relaxed success metrics to analyze baseline failure modes:
  • Distance-to-Goal-Only Success Rate (DTG-SR). Computed identically to HL-SR but without the semantic coverage requirement. This metric helps quantify instances where the agent selects frames close to the goal without necessarily capturing it visually.
  • Semantic Coverage Success Rate (SC-SR). Computed identically to HL-SR but without the distance-to-goal requirement. This metric evaluates instances where the agent correctly selects frames showing the target entity but fails to meet proximity criteria.
gemini-2.0-flash Task Prompt You are an intelligent question answering agent. In the prompt video preceeding the text, you will be shown a long stream of images that have been collected by a robot while navigating around a home and asked to answer a question about the space to assist a user in completing tasks in this home. In the video stream, the robot is picking up objects and rearranging them to different arbitrary locations throughout the house.
The questions asked by user will require being able to look at the full set of images collected by robot to be able to provide the answer. Your task is to identify the exact point in the video (by timestamp) where the user should move to best accomplish their goal. You can do this by outputting the exact timestep in the video where you are most confident that the object or place that they should move to, was viewed closely.
The images in the video also have the time of day information in top left. You need to use the time of day information for tasks that require to revisit objects or receptacles at a specific time of day.
For some tasks that require revisiting multiple receptacles or objects, you should output multiple frame indices corresponding to the correct order of revisitation (only if specified). For all user goals, identify the minimum number of objects or receptacles (targets) that you need to revisit to complete the user’s goal. For each revisitation, you should select one corresponding timestep in the video that you will move to. You need to ensure that you are not revisiting targets that are not relevant to the user’s goal.
Generate timestamps in the MM:SS format where the first two digits represent minutes and the last two digits represent seconds. To solve the tasks effectively, you should try to summarize the video by listing out the objects that were picked and placed, the receptacles from which each object was picked up and the receptacle where it was dropped.
The user’s goal is to: {goal}. Now look at the full video containing all the images. Give your final response with NUM_TARGETS_TO_REVISIT: ||<num_target>||. TIMESTAMP_INDEX: ||<timestamp_1>,…,<timestamp_num_target>||. You can include multiple timestamps if the task requires visiting more than one target (one timestamp for each target). Do not add any other text.
Figure 6: System Prompt for gemini-2.0-flash model
Appendix C Extended Related Work
C.1 An overview of approaches tackling memory in Long VQA and embodied tasks
Many recent works equip large language models with memory-like capabilities through task-specific knowledge bases or semantic maps for navigation planning [25, 8]. MobilityVLA [45] leverages long-context VLMs to process past frames, enabling goal-frame selection via Gemini, while topological mapping guides navigation without explicit pathfinding. ReMEmbR [3] and Embodied RAG [43] introduce nonparametric memory trees to store and retrieve past experiences for planning. For video question answering, Zhang et al. [49], Kahatapitiya et al. [13] propose a streaming-based approach that accumulates frame-wise captions for later querying. However, this method constrains memory to the expressivity of vision-language or captioning models, potentially omitting critical details. We include the hierarchical approach involving long-context VLMs and the approach utilizing textual memory traces as baselines in our evaluations on the benchmark tasks.
Appendix D Baseline Details
D.1 Gemini Agent
We experiment with the gemini-2.0-flash model which can process upto 1 million tokens in the context window. We strictly follow the outlined Google Gemini API developer guidelines [1] to structure the prompt and perform API calls over the long videos in the FindingDory tasks. We do not use a “structured output” format such as json in the output responses as we empirically observed it failed to produce coherent outputs when supplied with long videos in the context. Since the gemini-2.0-flash model expects the input frames to be sampled at 1 FPS, we prompt the model to directly generate specific timestamps in the recommended MM:SS format to localize the frame corresponding to the target entity. We provide the detailed prompt that we use for evaluations in Fig. 6.
D.2 Qwen and Gemma Agent
We run the Qwen2.5-VL-7B [5] and Gemma3-12B [40] locally for evaluations on the FindingDory tasks. The Qwen model uses a maximum of $768$ frames sampled uniformly across the input sequence while Gemma does not impose a maximum frame limit. For evaluation on 768 frames, we use two A40 GPUs with 48GB of memory. For experiments with shorter context lengths and higher subsampling (24–384 frames), a single A40 GPU suffices. We provide the full task prompt we use for inference with Qwen and Gemma in Fig. 7 and utilize structured JSON outputs.
Qwen/Qwen2.5-VL-7B-Instruct and google/gemma-3-12b-it Task Prompt You are an expert and intelligent question answering agent. You will be shown a video that was collected by a robot yesterday while navigating around a house and picking and placing objects. Your job is to help the robot complete a task today by looking at the video and finding the frame indices that the robot should move to. Make sure your response is all within the JSON.
Note:
- The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
Output Format (should be in JSON format):
‘‘‘json
  "chain_of_thought": str.
  "frame_indices": list[int].
Where:
- Chain of Thought: A detailed explanation of the robot’s thought process in determining the frame indices.
- Frame Indices: The frame indices in the video stream showing the object or place that the robot should move to.
- When deciding which frame indices to choose, make sure you choose the frame indices that are closest to the object/place.
- If task requires the agent to go to multiple places, output one frame index per object/place. (Do not use ellipsis)
The robot’s goal is: {goal}
Figure 7: System Prompt for Qwen2.5-VL and Gemma3 model
D.3 GPT Agent
We use the official OpenAI GPT-4o [12] API to run evaluations on the FindingDory tasks with video frames subsampled at $96$ frames to optimize API costs. For running complete evaluation on the entire validation task suite, we incurred a total cost of $\approx 400$ USD. We provide the full system prompt used for GPT-4o evaluations in Fig. 8.
gpt-4o-2024-07-18 Task Prompt ** OBJECTIVE **
You are an expert and intelligent question answering agent. You will be shown a video that was collected by a robot yesterday while navigating around a house and picking and placing objects.
Your job is to help the robot complete a task today by looking at the video and finding the frame indices that the robot should move to, to complete the task.
** OUTPUT FORMAT **
‘‘‘json
  "chain_of_thought": str.
  "frame_indices": list[int].
Where:
- Chain of Thought: A detailed explanation of the robot’s thought process in determining the frame indices.
- Frame Indices: The frame indices in the video stream showing the object or place that the robot should move to.
- When deciding which frame indices to choose, make sure you choose the frame indices that are closest to the object/place.
- If task requires the agent to go to multiple places, output one frame index per object/place.
** NOTES **
- The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
** TASK **
The robot’s goal is: {goal}
Figure 8: System Prompt for gpt-4o model
D.4 Text Agent
We implement the the text-based VLM agent using the Qwen2.5-VL [5] model. For every chunk_size number of frames in the video, we generate a description using the VLM. We then attach the frame indices and time of day information to the generated summary for the frame window in consideration. We use the prompt shown in Fig. 9 to generate the structured JSON outputs for each chunk in the video. We then concatenate all the generated JSON outputs and pass it to the Qwen model to perform text-only reasoning over the entire text-based history of the video. The text-based reasoning prompt is provided in Fig. 10. We note that using the same Qwen model for generating the individual text summaries and performing the final text-based reasoning helps in optimizing inference load as we maintain only a single instance of the VLM through the entire evaluation.
Text Agent(VLM) Task Prompt You are an expert and intelligent question answering agent. You will be shown a stream of {chunk_size} images taken by a robot while navigating around a simulated house and picking and placing objects. Your job is to describe the image based on the output format. Make sure your response is all within the JSON format.
Output Format (should be in JSON format):
‘‘‘json
  "room_name": str.
  "picking_placing_or_navigating": str("picking", "placing" or "navigating").
  "object_being_manipulated": str.
  "receptacle_being_manipulated": str.
  "other_objects_in_scene": list[str].
Where:
- Room name: The name of the room the robot is in.
- Picking, placing, or navigating: Whether the robot is picking or placing an object, or navigating around the house.
- Object being manipulated: The object that the robot is picking or placing (if relevant).
- Receptacle being manipulated: The receptacle that the robot is picking up from or placing an object into (if relevant).
- Other objects in scene: Objects present in the scene besides the main object and the receptacle.
Note:
- The robot uses a magic grasp action to pick up an object, where a gripper goes close to the object and the object gets magically picked up.
Figure 9: System Prompt for Qwen2.5-VL model to generate text summaries based on chunk_size frames
Text Agent(LLM) Task Prompt You are an expert and intelligent question answering agent. You will be provided a list of textual descriptions of the environment created by an agent while navigating around a house and picking and placing objects. The descriptions correspond to logs made after collecting and viewing a small chunk of frames, and contain information about which frame numbers they correspond to. You will also be provided a goal that now needs to be accomplished, for which you will need to use the history to decide where to go. Your job is to identify the desired frame index to navigate to, based on the provided task. Make sure your response is all within the JSON.
Agent’s History information:
- Room name: The name of the room the agent was in.
- Picking, placing, or navigating: Whether the robot was picking or placing an object, or navigating around the house.
- Object being manipulated: The object that the robot was picking or placing (if relevant).
- Receptacle being manipulated: The receptacle that the robot was picking up from or placing an object into (if relevant).
- Other objects in scene: Objects present in the scene besides the main object and the receptacle.
Output Format (should be in JSON format):
‘‘‘json
  "chain_of_thought": str.
  "frame_indices": list[int].
Where:
- Chain of Thought: A detailed explanation of your thought process in determining the frame indices.
- Frame Indices: A list of one or more frame indices relevant to accomplishing the goal.
- If its a single object/place task, output only one frame index.
- If task requires the agent to go to multiple places, output one frame index per object/place. (Do not use ellipsis)
Agent’s history: history
The goal is {goal}
Figure 10: System Prompt for Qwen2.5-VL model to perform textual reasoning over the concatenated textual summaries
D.5 Supervised Finetuning Baseline
We leverage the FindingDory training split which consists of $82174$ unique task instantiations across the categories mentioned in Table 2. Each video-instruction pair is coupled with a list of ground-truth frame indices that solve the task. For multi-goal tasks, each subgoal has a corresponding ground-truth frame indices sublist. Since the FindingDory video sequences can be extremely long, we uniformly subsample all videos to $96$ frames. The ground truth frame indices are also subsampled to match the shorter video lengths to ensure the model chooses from within the subsampled frames.
Training Details. We conduct full-finetuning of the Qwen/Qwen2.5-VL-3B model using the open-source huggingface/trl library. For hyperparameter optimization, we utilized a smaller representative split of the training dataset with $36274$ samples and conducted a grid search over the learning rate, num_epochs and weight_decay parameters. The best checkpoints from all runs were selected by evaluating performance on all video-instruction pairs in the full validation split. This “offline” evaluation is conducted without instantiating the simulator by directly comparing the model’s predicted keyframe indices with the ground-truth keyframe indices using exact string matching (thus, we dont compute metrics proposed in Section 4.2). We use the cosine learning rate decay scheduler for all experiments. For the final training on the larger dataset, we found that a learning rate of $5\times 10^{-6}$ over $5$ epochs with 00 weight decay worked best. For all experiments, we train in bfloat16 format using 8 A40 GPUs on a single node with a batch size of $1$ and gradient_accumulation of $4$ providing an effective batch_size of $32$ samples. To further optimize training, we use gradient_checkpointing and highly optimized flash_attention implementation. In this setup, the full training run takes $\approx 120$ hrs to complete.
Goal Sampling for Online Simulator Eval. The best trained checkpoint is used in online simulator evaluations on the FindingDory task suite. Since the model is trained to predict the complete list of ground-truth frame indices, we select a single representative index from each generated subgoal solution list to serve as the predicted frame index. Empirically, we found that choosing either the first or a random index (of a sublist corresponding to a subgoal) yields comparable performance, while using the last index results in slightly worse performance. Therefore, we use the first index of each list as a simple and effective heuristic.
Figure 11: Hierarchical agent architecture when using a VLM for goal frame selection and low level policies for action generation.
Appendix E Low-level Policy Details
E.1 Image-Goal Navigation Agent
Policy Training. We train the image-goal navigation policy using the architecture and training recipe described in [46]. Specifically, we employ an end-to-end reinforcement learning policy trained using DDPPO, which predicts discrete navigation actions conditioned on visual RGB inputs and a goal image. The agent utilizes a ViT based visual encoder along with a 2-layer LSTM backbone. Similar to OVRL-v2 [46], the output patch representations from the ViT are reshaped into a 3D grid and downsampled to a lower dimension using a convolutional layer called the compression layer. However, we slightly deviate from OVRL-v2 by first concatenating the patch representations from the current and goal images before passing them through the compression layer.
The policy training occurs in the training scene split of the HSSD dataset [15], consisting of $1166$ episodes distributed across $111$ scenes (a subset of the original $125$ training scenes). In the Habitat simulator, the agent is modeled as a Hello Robot Stretch [14] with a height of $1.41$ m and a cylindrical base radius of $0.25$ m. The agent’s RGB sensor is positioned at a height of $1.31$ m, with a resolution of $160\times 120$ pixels and a horizontal field of view of $43^{\circ}$. During training, each episode has a maximum step budget of $1000$ steps, with success defined by invoking the STOP action within $1.0$ m of the goal image position.
As our visual encoder, we use the VC-1-Base [22] model which was previously finetuned with ImageNav on a smaller training scene split from the HSSD dataset. Empirically, we found that freezing the visual encoder after this targeted finetuning significantly accelerates the overall training performance. During the VC-1 finetuning, we use a smaller learning rate of $1.5\times 10^{-6}$ for the encoder.
We train the agents for a total of $500$M timesteps on 32 A40 GPUs running 32 parallel environments each (1024 envs in total). Following the approach from [22], we collect $64$ steps of experience and subsequently perform 2 PPO epochs with a mini-batch size of $2$. The reward function employed is the improved formulation proposed by [46], using parameters: success weighting $c_{s}=5.0$, angle success weighting $c_{a}=5.0$, goal radius $r_{g}=1.0$, angle threshold $\theta_{g}=25^{\circ}$, and slack penalty $\gamma=-0.002$. We optimize using AdamW with a learning rate of $2.5\times 10^{-4}$ and weight decay $10^{-6}$.
We select the best-performing policy checkpoint based on a validation success rate obtained by evaluating on $1000$ image-goal navigation episodes from the FindingDory validation set, achieving a success rate of 78% and SPL of 48.21%.
E.2 Explicit Mapping Agent
We augment the OracleAgent (see Section B.2) with a top-down occupancy map as it navigates the environment to generate the video sequence. In addition, we also store the 2D poses corresponding to each frame that is collected in the video sequence. The constructed map is employed by a path planning module to generate a deterministic navigation sequence to the 2D map coordinates that correspond to the image-goal location selected by the high-level goal selection agent (see Section 4.1).
Map Construction. We leverage the onboard depth sensor and project the depth image at each timestep into an egocentric pointcloud using camera intrinsics. The pointcloud is then binned along the z-axis to compute occupancy values on a local 2D grid map centered around the current position of the agent. The pose sensor provides the instantaneous pose of the agent which is employed to register the computed local map in the global map of the environment. In this way, we continuously update the global map as the OracleAgent navigates the environment and visits multiple pick-place targets. We freeze the map updates during the pick-place subroutines as no additional occupancy information is encountered during their execution.
Planner. When the high-level goal agent selects a particular goal frame, we lookup the corresponding target 2D pose in the global map. Simlar to the local navigation policy employed in [8], we use the Fast Marching Method [34] to generate the shortest path to the goal pose. We use one of four discrete navigation actions (see Section 3.3) to reach each waypoint on the generated shortest path.
Appendix F Qualitative Examples
In this section, we include representative qualitative examples from the FindingDory task suite. In Figure 12, we provide a birds-eye view depiction of the trajectory executed during experience collection (see Section B.2) in a validation episode involving 5 pick-place interaction routines. For this episode, the collected video consists of $1308$ frames in total. We provide detailed responses (and meta-analysis) from various baselines across the various task categories in Figures 14, 15, 16, 17, 18 and 19. In the following, we briefly discuss the different failure modes encountered in the generated responses across the 3 high-level task categories defined in Table 2.
Single-Goal Spatial Tasks. We observe that for the Object Recall task (Fig. 14), the models are able to identify a frame with the target chest of drawers but do not select the closest frames to ensure success. In the Interaction based tasks (Figs. 14 and 15), we observe models fail to accurately list the sequence of interaction events which can be partially attributed to supplying only $96$ subsampled frames from the original video. In one case, GPT-4o instead focuses on a “distractor” target object (Fig. 15). All models perform poorly on the Spatial Relationship tasks (Fig. 15) as they are unable to coherently reason about the spatial layout of the house in the provided video.
Single-Goal Temporal Tasks. For the Interaction Order task, we observe that the closed-source models are able to accurately identify the target object but the open-source variants fail to do so (Fig. 17). Models are generally able to select the correct frame in the Time-Based task as the model only needs to retrieve a single frame corresponding to the correct overlaid timestamp. For the Duration Tracking task, only GPT-4o generates the correct reasoning and frame prediction. Intersetingly, Gemma-3 tries to track the duration spent in each room by matching the floor color observed in the images but eventually fails at the task (Fig. 18).
Multi-Goal Tasks. All models fail to perform successfully on tasks under this category. Only GPT-4o gets a correct response for all subgoals. Surprisingly, the reasoning trace has a hallucination but this mistake is omitted in the final frame sequence prediction (Fig. 19). The models struggle to coherently list the event sequence and also hallucinate interactions. Furthermore, Gemini-2.0-flash also struggles to adapt the sequence of events it lists into the query revisitation sequence specified in the task (Fig. 19).
Figure 12: An example experience collection trajectory executed by OracleAgent.
Figure 13: Example “magic” pick-place sequences executed by the OracleAgent for sample episode in Fig. 12.
Figure 14: FindingDory Spatial Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 15: FindingDory Spatial Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 16: FindingDory Spatial Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 17: FindingDory Temporal Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 18: FindingDory Temporal Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Figure 19: FindingDory Multi-Goal Tasks for the sample episode in Figure 12. Each VLM model uses $96$ subsampled frames. Text Agent processes full video as defined in Sec. D.4.
Table 3: Task Categories with all associated prompts
Memory Type Example Instructions
Navigate to a {target_object_name}.
Object Recall Navigate to a {target_receptacle_name}.
Navigate to any receptacle you interacted with.
Navigate to any receptacle you did not interact with.
Navigate to any object that you interacted with yesterday.
Navigate to any object that you did not interact with yesterday.
Navigate to any receptacle you picked an object from.
Interaction Navigate to any receptacle you placed an object from.
Navigate to the {target_object_name} that you interacted with yesterday.
Navigate to a {target_object_name} that you did not interact with yesterday.
Navigate to a {target_receptacle_name} you did not interact with yesterday.
Navigate to a {target_receptacle_name} you picked an object from.
Navigate to a {target_receptacle_name} you placed an object on.
Navigate to the receptacle that you picked the {target_object_name} from.
Conditional Interaction Navigate to the object that you picked from the {target_receptacle_name}.
Navigate back to a {target_shape} shaped object that you interacted with yesterday.
Navigate back to a {target_color} colored object that you interacted with yesterday.
Navigate to an interacted object with {target_print_or_design} on it.
Find an already interacted object that is made of {target_material}.
Object Attributes Go back to an interacted object that is used for {target_functionality}.
Navigate to the receptacle that you interacted with which is the farthest from your current location.
Navigate to the receptacle that you did not interact with which is the farthest from your current location.
Navigate to the receptacle that you picked an object from which is the farthest from your current location.
Navigate to the receptacle that you placed an object on which is the farthest from your current location.
Spatial Relationship Navigate to the object which you interacted with which is the farthest from your current location.
Navigate to the room where you picked the {target_appearance_order} object from.
Navigate to the room where you placed the {target_appearance_order} object in.
Navigate to the room where you picked the {target_object_name} from.
Navigate to the room where you placed the {target_object_name} in.
Room Visitation Navigate to a room that you did not visit yesterday.
Navigate to the {target_appearance_order} object that you interacted with yesterday.
Navigate to the {target_appearance_order} receptacle that you picked an object from.
Navigate to the {target_appearance_order} receptacle that you placed an object on.
Navigate to the receptacle that you picked the {target_appearance_order} object from.
Navigate to the object that you picked from the {target_appearance_order} receptacle.
Navigate to the object you interacted with immediately after ending the interaction with {target_object_name}.
Navigate to the object you interacted with immediately before interacting with {target_object_name}.
Navigate to the object you interacted with {target_num_counts} interactions after {target_object_name}.
Navigate to the object you interacted with {target_num_counts} interactions before {target_object_name}.
Navigate to the object that you interacted with between the interactions with {target_object_name_1} and {target_object_name_2}.
Navigate to the receptacle that you placed an object on right before you started interacting with {target_object_name}.
Navigate to the receptacle that you picked an object from right after you finished interacting with {target_object_name}.
Navigate to the receptacle that you placed an object on {target_num_counts} interactions before you started interacting with {target_object_name}.
Interaction Order Navigate to the receptacle that you picked an object from {target_num_counts} interactions after you finished interacting with {target_object_name}.
Navigate to the receptacle that you placed an object on between the interactions with {target_object_name_1} and {target_object_name_2}.
Navigate to the receptacle that you picked an object from between the interactions with {target_object_name_1} and {target_object_name_2}.
Navigate to the receptacle that you interacted with at {XX:XX} yesterday.
Time-Based Navigate to the object that you interacted with at {XX:XX} yesterday.
Navigate to the object which took the longest time to rearrange.
Navigate to the room that you spent the most time in.
Duration Tracking Navigate to the object which took the shortest time to rearrange.
Revisit all the receptacles you picked objects from yesterday.
Revisit all the receptacles you placed objects on yesterday.
Revisit all the {target_receptacle_name} you placed objects on yesterday.
Revisit all the {target_receptacle_name} you picked objects from yesterday.
Revisit all the objects you interacted with yesterday.
Unordered Revisitation Revisit all the receptacles you interacted with yesterday.
Revisit all the receptacles you picked objects from yesterday in specific order.
Revisit all the receptacles you placed objects on yesterday in specific order.
Ordered Revisitation Revisit all the objects you interacted with yesterday in specific order.